What is the solution of this system of equations?
-6y=-50-4x
5x-6y=-49

Answer by Alan3354(69443)

What is the solution of this system of equations?
-6y=-50-4x
5x-6y=-49
=================
-6y is solved for in terms of x.
Sub for -6y in the 2nd eqn.
5x + (-50-4x) = -49
Solve for x
etc